OpenStudy (anonymous):

Which of the following is a polynomial with roots 3, 5i, and −5i? A.) f(x) = x^3 - 3x^2 + 25x - 75 B.) f(x) = x^3 - 3x^2 + 15x - 25 C.) f(x) = x^3 - 15x^2 + 25x - 75 D.) f(x) = x^3 - 3x^2 + 15x - 75

OpenStudy (anonymous):

ax^3+bx^2+cx+d has three roots and imagine x,y and z are it's roots. then x+y+z=-b/a xyz=-d/a xy+zx+yz=c/a

OpenStudy (anonymous):

the roots are 5i,-5i,3 x+y+z=-b/a 3=-b/a b/a=-3 xyz=3*25=75=-d/a d/a=-75 in all of answers a=1so b=-3 and d=-75 so answer is x^3 - 3x^2 + 25x - 75